The Shmidt Valley Investment club invested $10000 at 5.1% annual simple interest. How much additional money should they invest at an annual simple interest rate of 3.5%, so that the total annual interest earned from the two accounts is 4.4% of the total investment.
Thanks, I tries this one a million times and no can do!

Answer by Paul(988)

51x+35(10000-x)=44(10000)
51x+350000-35x=440000
16x=90000
x=5625
hence, they should invest addition of $5625